OpenStudy (anonymous):

A jar contains 26 coins totaling $2.55 in nickels and dimes. How many dimes are in the jar?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

Let's say that x = number of nickles, and y = number of dimes. One nickle is $0.05 and one dime is $0.10. Establish two equations with x and y: x + y = 26 x*0.05 + y*0.10 = 2.55 From the first equation, let's find x: x = 26 - y Put this value into the second equation: (26 - y)*0.05 + y*0.10 = 2.55 To makes things easier we can multiply both sides of the equation by 100: (26 - y)*5 + y*10 = 255 -----> 130 - 5y + 10y = 255 -----> 5y = 125 -----> y = 125/5 -----> y = 25 So there are 25 dimes, which is your answer. To confirm the answer, notice again that x = 26 - y -----> x = 26 - 25 -----> x = 1 So there is only one nickle. We can confirm these answer by putting them into the second equation: x*0.05 + y*0.10 = 2.55 1*0.05 + 25*0.10 = 2.55 0.05 + 2.5 = 2.55 2.55 = 2.55

OpenStudy (lgbasallote):

let n = nickels d = dimes n+d = 26 5cn + 10cd = 255c (c means cent) systems... -5n - 5d = -130 5n + 10d = 255 5d = 125 d = 25 n = 1